# Dyonic Taub-NUT-AdS Black Branes: Thermodynamics and Phase Diagrams

### v1

@@ 3.2.2 Case II: $0<\absolutevalue{\alpha}\leq\frac{\sqrt{3}}{3}$
Case III.I: $\widetilde{Q}_{m}\in[-\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}},\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}]$
@@
Case III.II: $\widetilde{Q}_{m}\in[-\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6},-\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}})$ or $\widetilde{Q}_{m}\in(\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}},\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}]$
As stated earlier, the number of branches of the $T_{h}(r_{h})$ solution varies due to the value of $n$. In this case, there are four critical values of $n$,
$n_{c_{1}}=-\frac{3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}},$
$n_{c_{2}}=\frac{-3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}},$
$n_{c_{3}}=\frac{3\alpha^{2}l\phi_{e}^{2}-\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}},$
$n_{c_{4}}=\frac{3\alpha^{2}l\phi_{e}^{2}+\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}{3\,\alpha\,\phi_{e}}.$ (74)
@@
$r_{h(c_{2})}=\frac{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}+l\widetilde{Q}_{m}\right)-3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}}{3},$
$r_{h(c_{3})}=\frac{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)-3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}}{3},$
$r_{h(c_{4})}=\frac{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)+3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}}{3}.$ (75)
@@
$T_{c_{2}}=\frac{2}{3\pi l^{2}}\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}+l\widetilde{Q}_{m}\right)-3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}},$
$T_{c_{3}}=\frac{2}{3\pi l^{2}}\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)-3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}},$
$T_{c_{4}}=\frac{2}{3\pi l^{2}}\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)+3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}.$ (76)
(a) $-\infty<n<n_{c_{2}}$
@@
Figure 3.15: The temperature of the horizon of the black brane as a function of its horizon radius in case III.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=145$). Here we split the graph into three for clarity.
From Fig. 3.15 we can see that for $n\in[n_{c_{1}},n_{c_{2}}]$ or $n\in[n_{c_{3}},n_{c_{4}}]$ the horizon temperature of the black brane, $T_{h}(r_{h})$, has only one branch. For $n<n_{c_{1}}$ or $n>n_{c_{4}}$, it has three branches. Finally, for $n\in(n_{c_{2}},n_{c_{3}})$, it also has three branches except for $n=n_{hp}=\frac{\widetilde{Q}_{m}}{\alpha\phi_{e}}$ where $T_{h}(r_{h})$ has two branches and a minimum value at $r_{h}=r_{min}$ given by Eq.(73) which corresponds to $T_{min}$ given by Eq.(72).
@@
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, $\;n\in(n_{hp},n_{c_{3}})$, or $\;n>n_{c_{4}}\,$ and $\,T<T_{min}$ or $T>T_{max}\,$, the only stable solution is one black brane solution.
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, $\;n\in(n_{hp},n_{c_{3}})$, or $\;n>n_{c_{4}}\,$ and $T\in[T_{min},T_{tr})$, there are two stable black brane solutions, the small black brane is thermodynamically preferred while the big black brane is a metastable state.
@@
  • For $\,n<n_{c_{1}}$, $\;n\in(n_{c_{2}},n_{hp})$, $\;n\in(n_{hp},n_{c_{3}})$, or $\;n>n_{c_{4}}\,$ and $T\in(T_{tr},T_{max}]$, there are two stable black brane solutions, the big black brane is thermodynamically preferred while the small black brane is a metastable state.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$ or $n\in[n_{c_{3}},n_{c_{4}}]$, the only stable solution is one black brane solution.
  • For $n=n_{hp}$ and $T<T_{min}$, there are no possible solutions.
  • For $n=n_{hp}$ and $T\geq T_{min}$, the only stable solution is one black brane solution.
@@
Figure 3.18: The $n-T$ phase diagram for case III.II. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=145$). The gray line at $n=n_{hp}$ shows the region where no black brane can exist.
@@
Figure 3.19: The $n-T$ phase diagram for case III.II where two of the critical points merge. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=150$) The gray line at $n=n_{hp}$ shows the region where no black brane can exist.
Case III.III: |$\absolutevalue{\widetilde{Q}_{m}}>\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}$
As the value of |$\absolutevalue{\widetilde{Q}_{m}}$ | increases to be greater than $\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}$ the two merged critical points vanish completely and the case becomes exactly the same as Case II.III.
@@ 4 Conclusion
We started with a review of the phase structure of the NUT-less dyonic black branes in the mixed ensemble, i.e. an ensemble with fixed electric potential and magnetic charge. Also, we have calculated the thermodynamic quantities of the solution and show that they obey the first law and Gibbs-Duhem relation. We showed that in the NUT-less case the phase structure is trivial since there are no phase transitions and the temperature is a monotonic function of the black hole horizon radius.
In the main work of the paper we have applied the thermodynamic treatment introduced in [33, 34] to the dyonic-NUT-AdS spacetimes with flat horizons to obtain various thermodynamic quantities. We showed that these quantities satisfy the first law as well as the Gibbs-Duhem relation. It was intriguing to notice that the first law is satisfied by a larger class of charges and potentials which depends on some additional arbitrary parameter $\alpha$ which in turn affects the phase structure. We analyzed the phase structure of these solutions through a mixed ensemble where we described the phase diagrams as NUT parameter-temperature graphs to show possible transitions between big and small black hole phases. We showed that due to the absence of Misner strings, one doesn’t need to fix $\alpha$, therefore, it takes any value [35]. Depending on the value of $\alpha$, one can classify the phase study into different distinct cases which could be further split into few sub-cases depending on the value of the other thermodynamic quantities. In most of these cases we have first-order phase transitions which end at one or more critical points. We found that these classes could have up to four critical points, again, depending on $\alpha$. It is important to notice that these phase structures are qualitatively different from the van der Waals structure of the liquid-gas phase transition.

### v2

@@ 3.2.1 Case I: $\alpha=0$
[…]
@@
[…]
As stated earlier, the monotonicity of $T_{h}(r_{h})$ varies due to the value of $n$. In this case, we have all the four critical points
$n_{c_{1,2}}=-\alpha l\phi_{e}\mp\frac{1}{3}{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}},$ (83)
@@
$r_{h(c_{1,2})}=\frac{1}{3}{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}+l\widetilde{Q}_{m}\right)\pm 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}}}\,$ (84)
@@
$T_{c_{1,2}}=\frac{2}{3\pi l^{2}}\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}+l\widetilde{Q}_{m}\right)\pm 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)+6\widetilde{Q}_{m}\right)}},$ (85)
and
$n_{c_{3,4}}=\alpha l\phi_{e}\mp\frac{1}{3}{\sqrt{3l\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}},$ (86)
corresponding to
$r_{h(c_{3,4})}=\frac{1}{3}{\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)\mp 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}}},$ (87)
and
$T_{c_{3,4}}=\frac{2}{3\pi l^{2}}\sqrt{9\left(\alpha^{2}l^{2}\phi_{e}^{2}-l\widetilde{Q}_{m}\right)\mp 3l\sqrt{3\alpha^{2}l\phi_{e}^{2}\left(l\phi_{e}^{2}\,(3\alpha^{2}+1)-6\widetilde{Q}_{m}\right)}},$ (88)
(a) $\infty<n<n_{c_{2}}$
@@
Figure 3.15: The black brane horizon temperature as a function of the horizon radius in case III.II for different values of $n$. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=145$). Here we split the graph into three for clarity.
From Fig. 3.15 we can see that for $n\in[n_{c_{1}},n_{c_{2}}]$ or $n\in[n_{c_{3}},n_{c_{4}}]$ the horizon temperature of the black brane, $T_{h}(r_{h})$, is monotonic. For $n<n_{c_{1}}$ or $n>n_{c_{4}}$, $T_{h}(r_{h})$ is non-monotonic, having three branches. Finally, for $n\in(n_{c_{2}},n_{c_{3}})$, it also has three branches except for $n=n_{hp}=\frac{\widetilde{Q}_{m}}{\alpha\phi_{e}}$ where $T_{h}(r_{h})$ has two branches and a minimum value $T_{min}$ given by Eq.(81) at $r_{h}=r_{min}$ given by Eq.(82).
@@
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, $n\in(n_{hp},n_{c_{3}})$, or $n>n_{c_{4}}\,$ and $T<T_{min}$ or $T>T_{max}\,$, the only stable phase is one black brane.
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, $n\in(n_{hp},n_{c_{3}})$, or $n>n_{c_{4}}\,$ and $T\in[T_{min},T_{tr})$, there are two stable black brane phases, the small black brane is thermodynamically preferred while the big black brane is metastable.
@@
  • For $n<n_{c_{1}}$, $n\in(n_{c_{2}},n_{hp})$, $n\in(n_{hp},n_{c_{3}})$, or $n>n_{c_{4}}\,$ and $T\in(T_{tr},T_{max}]$, there are two stable black brane phases, the big black brane is thermodynamically preferred while the small black brane is metastable.
@@
  • For $n\in[n_{c_{1}},n_{c_{2}}]$ or $n\in[n_{c_{3}},n_{c_{4}}]$, the only stable phase is one black brane.
  • For $n=n_{hp}$ and $T<T_{min}$, there are no possible black brane phases.
  • For $n=n_{hp}$ and $T\geq T_{min}$, the only stable phases is one black brane.
@@
Figure 3.18: The nTn-T phase diagram for case III.II. The values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=145$). The gray line at $n=n_{hp}$ shows the region where black brane cannot exist as a phase.
@@
Figure 3.19: The nTn-T phase diagram for case III.II where two of the critical points merge. Values used to draw this graph are ($l=1$, $\phi_{e}=15$, $\alpha=1$, $\widetilde{Q}_{m}=150$) The gray line at $n=n_{hp}$ shows the region where black brane cannot exist as a phase.
[…]
As the value of $|\absolutevalue{\widetilde{Q}_{m}}$ increases to be greater than $\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}$ the two merged critical points vanish completely and the case becomes exactly the same as Case II.III.
@@ 4 Conclusion
We started with a review of the phase structure of the NUT-less dyonic black branes in the mixed ensemble, i.e. an ensemble with fixed electric potential and magnetic charge. Also, we have calculated the thermodynamic quantities of the solution and show that they obey the first law and Gibbs-Duhem relation. We showed that in the NUT-less case the phase structure is trivial since there are no phase transitions and the temperature is a monotonic function of the black brane horizon radius.
In the main work of the paper we have applied the thermodynamic treatment introduced in [5, 6] to the dyonic-NUT-AdS spacetimes with flat horizons to obtain various thermodynamic quantities. We showed that these quantities satisfy the first law as well as the Gibbs-Duhem relation. It was intriguing to notice that the first law is satisfied by a larger class of charges and potentials which depends on some additional arbitrary parameter $\alpha$ which in turn affects the phase structure. We analyzed the phase structure of these solutions through a mixed ensemble where we described the phase diagrams as NUT parameter-temperature graphs to show possible transitions between big and small black brane phases. We showed that due to the absence of Misner strings, one doesn’t need to fix $\alpha$, therefore, it takes any value [51]. Depending on the value of $\alpha$, one can classify the phase study into different distinct cases which could be further split into few sub-cases depending on the value of the other thermodynamic quantities. In most of these cases we have first-order phase transitions which end at one or more critical points. We found that these classes could have up to four critical points, again, depending on $\alpha$. These cases are summarized in Table 1.
Range of $|\absolutevalue{\alpha}$ Range of $|\absolutevalue{\widetilde{Q}_{m}}$ Number of Critical Points
=0=0 =0=0 0
$\neq 0$ 2
$\in(0,\frac{\sqrt{3}}{3}]$ $\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$ 2
$\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$ 1
$\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}$ 2
$\frac{\sqrt{3}}{3}$ $\in[0,\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}}]$ 2
$\in(\frac{\alpha\,l\,\phi_{e}^{2}}{\sqrt{3}},\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}]$ 4
$\frac{l\phi_{e}^{2}\left(1+3\alpha^{2}\right)}{6}$ 2
Table 1: Number of thermodynamic critical points for the dyonic Taub–NUT–AdS black brane under different parameter ranges of $\alpha$ and $\widetilde{Q}_{m}$, corresponding to the sub-cases analyzed in Section 3.
It is important to notice that these phase structures are qualitatively different from the van der Waals structure of the liquid-gas phase transition.